OpenStudy (anonymous):

Find the distance between (3, 8) and (7, -11). Round to the nearest hundredth. I ended up getting the answer 20. Not sure if it's right...

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

I forgot it. :(

OpenStudy (anonymous):

\[d= \sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1}^2)}\]

OpenStudy (anonymous):

So now simply plug in the values into the formula.

OpenStudy (anonymous):

okay, thanks, I should be fine now that i have the formula :3

OpenStudy (anonymous):

Let me know what you get so we can check it :b

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

the answer should be ~19.42

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

is that a negative? :/

OpenStudy (anonymous):

no, positive! :o

OpenStudy (anonymous):

oh. but how'd you get 19.24?

OpenStudy (anonymous):

or 42 sorry

OpenStudy (anonymous):

ok, well lets do it one step at a time.. Find the distance between (3, 8) and (7, -11). Round to the nearest hundredth. =sqrt((x2-x1)^2+(y2-y1)^2) =sqrt((7-3)^2+(-11-8)^2) =sqrt(16+361) =sqrt(377) =19.4164878389476 =19.42 ((nearest hundredth))

OpenStudy (anonymous):

I substituted the values of the x and y coordinates of the 1st and 2nd set respectively

OpenStudy (anonymous):

i am.. completely lost at what you did right there..

OpenStudy (anonymous):

well i mean i know where you substituted the x and ys.. but

OpenStudy (anonymous):

then i set them to cubed.. i got the 16. But how'd you get the 361? i had gotten a 4. then put it cubed..

OpenStudy (anonymous):

you're supposed to square it, I guess is what you did.. Alright. I understand it a bit now.. I'll try that.

OpenStudy (anonymous):

=sqrt((x2-x1)^2+(y2-y1)^2) =sqrt((7-3)^2+(-11-8)^2) =sqrt(16+361) =sqrt(377) =19.4164878389476 =19.42 ((nearest hundredth)) well for (7-3)^2 = 4^2 the answer is 16. but for the y^2 values you are doing.. (-11-8)^2 = (-19)^2 = 19 squared is equal to 361.

OpenStudy (anonymous):

but wouldn't you have to turn the 8 into a negative?

OpenStudy (anonymous):

the 8 is already negative. subtract 8 from -11. = -19

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

:/ try solving it step by step, and ill check each one to find the error.